# Pathspace Kalman Filters with Dynamic Process Uncertainty for Analyzing Time-course Data

### v1

@@ 6 Related Work
Given their long history, numerous extensions and applications of the KF exist. Famously, KFs have been used in economics [4], parameter estimation, such as estimating insulin sensitivity [2], estimation of electro-chemical charge parameters [19], etc. Here we discuss some general extensions of the KF and their specific use in biology-related applications.
Extensions of the Kalman Filter: Simple extensions include multivariate KFs [5], non-linear or extended KFs [7], unscented KFs [8], and ensembles of KFs [13]. Of particular relevance to our work are Adaptive KFs as derived by Brown and Rutan, which dynamically adapts the filter’s measurement model to the observed data by matching the covariances in the internal model predictions with the covariances of the measurements [3]. Another inspiration for our work, is by Husain et. al., which demonstrates through both theoretical analyses and experiments in photosynthetic cyanobacteria a process of self-tuning the bio-physicial sensitivity to environmental cues, which can be understood within this filtering framework [34].
Recently, there has also been considerable work linking KFs to various machine learning approaches well summarized in a recent review [38]. Several notable examples including using neural networks to fit KF’s process uncertainty and data variance parameters [45], using a neural network as the KF’s internal model [27], and replacing the Gaussian representation of distributions in the KF with normalizing flow neural networks [35], and in hybrid learning models [39].
@@
Methods to Analyze Time-course Biological Data: Numerous methods have been developed to analyze time-course biological datasets. When using gene expression data, differential expression analysis is commonly employed [24]. However, this method does not explicitly take dynamics into account. Several methods have also employed KFs to analyze gene expression time-courses; [28] and [32] use KFs to model gene regulatory networks by detecting statistically significant changes in time-varying linear KFs and using KFs to construct a Bayesian network, respectively. A related approach is to fit a mechanistic model to the dynamical datasets which can be done with non-linear KFs [21] or other methods [40]. However, such approaches typically do not scale to large mechanistic models and suffer from issues related to parameter identifiability [47].
Finally, many methods fit univariate gene expression models RNAseq data [33] and in some cases use dynamically changing fit parameters to model temporal dynamics [41]. However, these models do not directly provide biological insight; instead the task of statistically analyzing the gene expression data is transformed into the related task of statistically analyzing the fit parameters.
@@ 7 Conclusion
In summary, our model provides a powerful extension of traditional KFs optimized for time-course datasets with the ability to dynamically highlight temporal regions where the model and data disagree.
Due to the model-driven nature of this work, every new domain requires its own custom internal model development. Although black-box Bayesian methods can be used, these will not scale well and for large data applications custom models must be developed. Currently, we have presented internal models for population dynamics and gene expression. In the future, we hope to add internal models for other biological experiments including single cell sequencing, DNA methylation, CHIPseq, and ATACseq measurements.
In the future, it may be possible to replace custom mechanistic models with black-box AI models. However, one advantage of the domain-specific models is that they require relatively few measurements to run effectively. A fully black-box approach would likely require much larger datasets, which are frequently not available in biological applications.
Finally, we comment that this work only deals with the univariate case which could be extended to a multivariate formalism. Similarly, the assumption of normality of all underlying distributions could potentially be relaxed in future extensions of this method.

### v2

@@ 6 Related Work
Table 3 and Figure 5 compare several related methods to the PKF which outperforms all of them on MSE. We discuss these methods below. Additionally Table B.1 in Appendix B compares the various similarities and differences between the benchmarked algorithms.
Extensions of the Kalman Filter: Simple extensions include multivariate KFs [multivariable_control], non-linear or extended KFs [extendedkf], unscented KFs [unscentedkf], and ensembles of KFs [ensemblekf]. Of particular relevance to our work are Adaptive KFs which dynamically adapts the filter’s measurement model to the observed data by matching the covariances in the internal model predictions with the covariances of the measurements [BrownandRutan]. Another inspiration for our work comes from analysis of the cyanobacteria circadian clock which is shown to be reminiscent of a Kalman filter non-monotonic Kalman gain [Husain19]. Finally, KFs have been linked to various machine learning approaches[Kim2022]. Several notable examples including using neural networks (NNs) to fit KF’s process uncertainty and data variance parameters [Greenberg2023], using a NN as the KF’s internal model [Wu2012], and replacing the Gaussian representation of distributions in the KF with normalizing flow NNs [deBezenac2020], and in hybrid learning models [levine2022framework]. The PKF draws inspiration from many of these methods - particularly with regards to estimating process uncertainty dynamically - but is more closely related to Bayesian smoothing algorithms in its construction.
Algorithm Parameters MSE
Adaptive Non-linear KF Q=1 125.56
Q=10 71.09
Unscented KF Q=1 50.13
Q=10 19.28
Unscented RTS Q=1 200.85
Q=10 131.56
IPLS Q=1, Iterations = 1 200.38
Q=1, iterations=10 132.3
Q=10, Iterations=1 121.17
Q=10, Iterations=10 11.38
PKF Iterations=1 4.36
Iterations=10 0.88
Table 3: Mean-square-error (MSE) of several methods on synthetic population dynamics data
Comparison to Bayesian Smoothing: Bayesian smoothing algorithms are a class of smoothing algorithms with roots in Gaussian processes. These algorithms frequently use KFs ability to account for different noise sources in order to smooth data. The RTS smoother and unscented RTS smoother both use KFs to estimate states looking forwards and backwards in time across an entire trajectory [RTS_smoother, UnscentedRTS]. The iterative posterior linearized smoother (IPLS) is an extention of the unscented RTS smoother which, similarly to a PKF, iterates the smoother output back into itself [IPLS]. However, none of these methods dynamically update process uncertainty and, as we show in Table (3), this parameter can have dramatic effects on the mean-squared-error. The PKF solves this problem by automatically updating its process uncertainty as it iterates.
Figure 5: Outputs of the methods benchmarked in Table 3.
Comparison to Parameter Estimation Algorithms: Parameter estimation is an immense topic, so here we focus on parameter estimation as it relates to KFs. Adaptive KFs attempt to estimate their internal parameters from data [brown1985adaptive]. Many extensions of KFs have also been developed to estimate model parameters as observable state variables [sarkka2023bayesian]. The PKF builds on these ideas by adaptively updating both its process uncertainty (via update equation 8) and the internal model parameters via the internal Bayesian model computation described in section 4. Unlike other approaches, the PKF uses analytically tractable internal models to directly compute Bayesian parameter posteriors which results in higher accuracy and more scalability. To our knowledge, no other methods iteratively update these parameters the way the PKF does.
Comparison to Change Point Detection Methods: Change point detection methods are a large subject, so here we focus on methods that make use of KFs. Threshold based change point detection algorithms have been used to determine when to re-compute internal KF parameters [PODDAR2016410_changepoint1]. Similarly to the PKF, the error of between filter estimates and data has been tracked to identify changes in the underlying model [changepoint2]. Finally, ensembles of KFs have been combined with a Naive Bayes’ algorithm for change point detection [Daumer1998OnlineCD_changepoint3]. The dynamic process uncertainty of the PKF accomplishes a similar goal to these approaches but without the use of any hyper parameters such as thresholds.
@@
Methods to Analyze Time-course Biological Data: When using gene expression data, differential expression analysis is commonly employed [anders2010differential]. However, this method does not explicitly take dynamics into account. KFs have been used to model gene regulatory networks by detecting statistically significant changes in time-varying linear KFs [Xiong2013] and using KFs to construct a Bayesian network [Pirgazi2018]. A related approach is to fit a mechanistic model to the dynamical datasets which can be done with non-linear KFs [Qian2008] or other methods [Linden2022]. However, such approaches typically do not scale to large mechanistic models and suffer from issues related to parameter identifiability [Lederer2024].
@@ 7 Conclusion
In summary, our model provides a powerful extension of KFs optimized for time-course datasets with the ability to dynamically highlight temporal regions where the model and data disagree. This method has been designed for use on biological datasets where identifying unexpected dynamics offers an opportunity for new scientific discovery. Towards this end, we have developed highly optimized Bayesian models for population dynamics and gene expression dynamics. We have applied these methods to real and synthetic datasets to demonstrate the accuracy and scalability of our approach.
Limitations: First, we comment that the PKF only deals with the univariate systems which could be extended to a multivariate formalism. Similarly, the assumption of normality of all underlying distributions could potentially be relaxed in future extensions of this method.
Second, fast parameter inference algorithms are necessary for our method to work on large datasets. We have presented a technique to implement these in cases where the internal model has a simple analytic solution. However, scaling up to large models remains an important challenge. Theoretically, black box Bayesian parameter inference methods could be used, but these will not scale well for large data applications such as those required in biology.
Although the PKF is undoubtedly more computationally intensive than some other filtering techniques, it is also more powerful and accurate. However, it is still fast enough to run on fairly large datasets without specialized compute resources and is therefore appropriate for many kinds of big data analyses. That said, for real-time data processing the PKF may not be appropriate.